OpenStudy (anonymous):

A 12-foot ladder is leaning against a wall. The bottom of the ladder is 5 feet away from the bottom of the wall. Approximately how high up the wall does the top of the ladder reach? A. 2.4 feet B. 7.0 feet C. 13.0 feet D. 10.9 feet

OpenStudy (anonymous):

Is it 10.9 feet?

OpenStudy (anonymous):

@Data_LG2 are you good with math??

OpenStudy (anonymous):

if you are can you help me?

rishavraj (rishavraj):

@Hermitberg yup its 10.9 .....

OpenStudy (anonymous):

ok tahnk you! :D

OpenStudy (anonymous):

can you help me with some others?

rishavraj (rishavraj):

hmmmm i shall try....

OpenStudy (anonymous):

A tree is 132 inches tall. The distance from the top of the tree to the tip of the tree's shadow is 166 inches. Approximately how long is the tree's shadow? A. 212 inches B. 149 inches C. 101 inches D. 34 inches http://static.k12.com/calms_media/media/1422500_1423000/1422752/1/1027944e04f03ab3880b4c7164012303b3883749/MS_PA_131002_171201.jpg

OpenStudy (anonymous):

is it A. ??

OpenStudy (anonymous):

if you don't want to click on the link I can try to draw it

rishavraj (rishavraj):

nah its not A....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

maybe it is B. ??

rishavraj (rishavraj):

oooo no its not tht...... http://prntscr.com/7hvckj

OpenStudy (anonymous):

oh its C. 101 inches

rishavraj (rishavraj):

yeah it is....

OpenStudy (anonymous):

cool thanks

OpenStudy (anonymous):

The correct answer is D. You use Pythagorean identity \( x^2 + y^2 =c^2 \) c = 12 x = 5 \( 5^2 + y^2 =12^2 \) \( y^2 =12^2 - 5^2 \) \( \sqrt{y^2} =\sqrt{12^2 - 5^2} = \sqrt{119} = 10.9 \)

OpenStudy (anonymous):

Just wanted to show the work just in case someone searches the question and see this :-)